OpenStudy (anonymous):

use f(x)=x+3 and g(x)=x+9 to find each of the following (a).(f+g)(x) (b).(f-g)(x) (c).(fg)(x) (d).(f/g)(x)

OpenStudy (anonymous):

a) x+3+x+9=2x+12 c) (x+3)(x+9)=x^2+12x+27 Try b and d yourself.

OpenStudy (anonymous):

K thank u....